OpenStudy (anonymous):

A bag contains 24 blue marbles, 32 yellow marbles, 43 white marbles, and 21 red marbles. If a marble is drawn from the bag and replaced 90 times, about how many times would a yellow marble be drawn?

jimthompson5910 (jim_thompson5910):

How many marbles are there total?

OpenStudy (anonymous):

120

jimthompson5910 (jim_thompson5910):

what fraction of that 120 are just yellow?

OpenStudy (anonymous):

32/120

jimthompson5910 (jim_thompson5910):

which reduces to ?

OpenStudy (anonymous):

4/15 ?

jimthompson5910 (jim_thompson5910):

That's the probability of drawing a yellow marble

jimthompson5910 (jim_thompson5910):

Multiply that probability by the number of trials (90) to get your final answer

OpenStudy (anonymous):

So I would have to convert the fraction to a decimal first right?

jimthompson5910 (jim_thompson5910):

or you can multiply the fraction with 90 think of 90 as 90/1

jimthompson5910 (jim_thompson5910):

then multiply straight across and reduce

OpenStudy (anonymous):

\[\frac{ 4 }{ 15} \frac{ 90 }{ 1 } 4\times90=360|15\times1=15| \frac{ 360 }{ 15 }?\]

OpenStudy (anonymous):

\[\frac{ 360 }{ 15 }\] is an improper fraction though?

jimthompson5910 (jim_thompson5910):

now reduce 360/15

OpenStudy (anonymous):

\[\frac{ 24 }{ 1 }\]

jimthompson5910 (jim_thompson5910):

and that's the same as 24

jimthompson5910 (jim_thompson5910):

so you'd expect about 24 marbles to be yellow (out of the 90)

jimthompson5910 (jim_thompson5910):

and notice how 24/90 reduces to 4/15 which is the probability of choosing yellow

OpenStudy (anonymous):

Cool :D Thanks +Medal/Fan

jimthompson5910 (jim_thompson5910):

you're welcome